# Are Vision xLSTM Embedded $U$-Nets More Reliable in Medical 3D Image Segmentation?

### v2

@@ 4.3 Results
Fig. 7 provides a graphical analysis of the parameter count, Tera FLoating point OPerations (TFLOPs), and memory size on disk, of our U-VixLSTM compared to that of other baseline architectures under consideration. The analysis shows that U-VixLSTM has the lowest number of parameters and TFLOPs (floating point operations per second) compared to the other SOTA. This reiterates our claim about the superior computational efficiency of the proposed model. Hence, it demonstrates potential for deployment in resource-constrained environments.
5 Conclusions and Discussion
@@
This research introduced the first integration of CNNs with Vision-xLSTMs for image segmentation (specifically, medical image segmentation), using the widely used $U$-shaped framework. CNNs acquired intricate spatial and textural information from the input image. The Vision-xLSTMs captured both global and temporal relationships inside the patches, derived from the feature volume produced by CNNs; thus learning a robust representation for the target structures.
@@
The superiority of our model U-VixLSTM can be attributed to various factors. The incorporation of CNNs with Vision-xLSTM allowed the model to effectively capture both local and global contextual information, which was essential to achieve precise segmentation. Furthermore, the utilization of Vision-xLSTM enhanced the efficiency of the model in terms of the total number of parameters and FLOPs, compared to other hybrid CNN transformer methods. This has substantial implications for clinical practice by increasing the accessibility of our model in a scenario with limited resources. This can potentially lead to an expedited and effective identification of medical conditions, decision-making for treatment, and providing guidance during surgical procedures.
@@ Acknowledgement
This work was supported by the J. C. Bose National Fellowship, grant no. JCB/2020/000033 of S. Mitra.

### v3

@@ 3.3 Results and Discussion
Fig. 11 provides a graphical analysis of the parameter count, Tera FLoating point OPerations (TFLOPs), and model size on disk, of our U-VixLSTM compared to that of other baseline architectures under consideration. The analysis shows that U-VixLSTM has the lowest number of parameters and TFLOPs (floating point operations per second) compared to the other SOTA. This reiterates our claim about the superior computational efficiency of the proposed model. Hence, it demonstrates potential for deployment in resource-constrained environments.
4 Conclusions
@@
This research introduced the first integration of CNNs with ViL for image segmentation (specifically medical image segmentation), using the widely used $U$-shaped framework. CNNs acquired intricate spatial and textural information from the input image. The ViL captured both global and temporal relationships inside the patches, derived from the feature volume produced by CNNs; thus learning a robust representation for the target structures.
@@
The superiority of U-VixLSTM can be attributed to various factors. Incorporation of CNNs with ViL allowed the model to effectively capture both local and global contextual information, which was essential to achieve precise segmentation. The exponential gating mechanism enables the model to efficiently retain longer contextual information necessary to precisely delineate target structures of variable sizes. Furthermore, the utilization of ViL improved the efficiency of the model in terms of the total number of parameters and FLOPs, compared to other hybrid CNN transformer methods. This has substantial implications for clinical practice by increasing the accessibility of our model in a scenario with limited resources. This can potentially lead to an expedited and effective identification of medical conditions, decision-making for treatment, and providing guidance during surgical procedures. Furthermore, the remarkable performance highlights the potential of ViL for effective medical image segmentation. This study lays the groundwork for further research and exploration in the application of ViL to this field.
@@ Acknowledgement
This research was supported by J. C. Bose National Fellowship grant JCB/2020/000033 of S. Mitra.